OpenStudy (anonymous):

Three wires are 6.5 m, 8.19 m, and 4.457 m long. What is the total length of the wires? Give your answer with the appropriate precision. A. 19 m B. 19.0 m C. 19.1 m D. 19.147 m

OpenStudy (anonymous):

What is 6.5+8.19+4.457 = .... ?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

Should an estimate or the exact answer be calculated? Then do the appropriate calculation. A charity raised money for a 6 week period. The group had a goal of collecting at least $500. They collected $69.75 in each of the first two weeks and $74.35 in each of the last four weeks. Did the charity reach it's goal? A. The exact answer is needed to know if the goal was reached. $436.90 was collected. B. An estimate can be used to know if the goal was reached. 70• 2 + 75 • 4 = 440 At most $440 was collected.

OpenStudy (anonymous):

We have : (2 * 69.75)+(4*74.35)=The money they can collecting . So if you reduce the money they want (500$)from that can answer .

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@skinny23 plzz help

OpenStudy (anonymous):

it is a because yes u need to add them all together and multiply them by how many weeks

OpenStudy (anonymous):

It's A anyway, wanted to teach you how to do it.

OpenStudy (anonymous):

:O ic sry plz go on

OpenStudy (anonymous):

yes drew is correct

OpenStudy (anonymous):

Should an estimate or the exact answer be calculated? Then do the appropriate calculation. Megan had a $50 gift card for the sporting goods shop. She made 3 purchases of $6.75 each and 2 purchases of $10.50 each. How much money is left on the card? A. An estimate is sufficient. There is about $9.00 left on the card. B. The exact answer is needed. There is $8.75 left on the card.

OpenStudy (anonymous):

Answer is B. Sorry for previous post

OpenStudy (anonymous):

np

OpenStudy (anonymous):

An electrician has a spool of wire 200 m long. She needs four pieces of wire measuring 42.3 m, 37.8 m, 55.6 m, and 35.7 m. How much wire will be left on the spool after she finishes her job? A. 20 m B. 28 m C. 28.6 m D. 29.6 m

OpenStudy (anonymous):

ok ic now

OpenStudy (anonymous):

Well, to figure this one out we need to do some subtraction: 200 - 42.3 = 157.7 , 157.7 - 37.8 = 119.9 , 119.9 - 55.6 = 64.3 , 64.3 - 35.7 = ???

OpenStudy (anonymous):

Oh, got it already?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im working on it lol

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Is it 28.6?

OpenStudy (anonymous):

Yes, good job!